OpenStudy (anonymous):

Please help solve! (x+1)/3 + (x+2)/7 = 5

OpenStudy (anonymous):

\[\frac{x+1}{3}+\frac{x+2}{7}=5\]

OpenStudy (anonymous):

you need to get common denominators

OpenStudy (anonymous):

how would i do that? I'm rusty :p

OpenStudy (johnweldon1993):

You need to find what number 3 and 7 both go into evenly

OpenStudy (anonymous):

I know its 21

OpenStudy (anonymous):

but what is that trick where you multiply the bottom by something and add it to the top? thats what i can't remember

OpenStudy (johnweldon1993):

So lets just say for the sake of anything ...that number is 21 (because it is :P) We need to multiply the whole first term by 7/7 (to make the denominator 21) and multiply the whole second term by 3/3 (to make that denominator 21 as well) \[\large \frac{7}{7}(\frac{x + 1}{3}) + \frac{3}{3}(\frac{x + 2}{7}) = 5\] \[\large \frac{7(x + 1)}{21} + \frac{3(x + 2)}{21} = 5\] Put both the terms over that common 21 \[\large \frac{7(x + 1) + 3(x + 2)}{21} = 5\] Multiply both sides by 21 \[\large 7(x + 1) + 3(x + 2) = 105\] Distribute as needed \[\large 7x + 7 + 3x + 6 = 105\] Combine like terms \[\large 10x + 13 = 105\] Subtract 13 from both sides of this equation \[\large 10x = 92\] Divide both sides by 10 \[\large x = 9.2\] Now lets check it \[\large \frac{x + 1}{3} + \frac{x + 2}{7} = 5\] becomes \[\large \frac{9.2 + 1}{3} + \frac{9.2 + 2}{7} = 5\] \[\large \frac{10.2}{3} + \frac{11.2}{7} = 5\] \[\large 3.4 + 1.6 = 5\] \[\large 5 = 5 \color \red{\checkmark} \]

OpenStudy (johnweldon1993):

That trick you're talking about would be when you multiply both the top AND bottom by the number needed to make the common denominator :)

OpenStudy (anonymous):

ohhh thank you SO much!! :)